OpenStudy (anonymous):

An earthworm of length 15cm is crawling along at 2cm/s . An ant overtakes the worm in 5 seconds how fast is the ant walking ? Many ppl explained this but I just dont get it :@

OpenStudy (anonymous):

An ant needs to travel at 3cm/s in order to bridge 15cm in 5 seconds. \[\frac{ 15 }{ 5 } = 3\] If a worm is crawling along at 2cm/s, the ant needs to travel 2cm/s faster than 5cm/s. \[5cm/s + 2cm/s = 7cm/s\] The ant is walking at 7cm/s.

OpenStudy (anonymous):

Are you sure , cause many ppl have answered it as 5cm ?

OpenStudy (anonymous):

Yes, I'm sorry, I meant \[3cm/s + 2cm/s = 5cm/s\]

OpenStudy (anonymous):

Thanks btw